OpenStudy (anonymous):

Use the Distance Formula to find the distance to the nearest tenth between A(-7, -4) and B(-2, 0).

OpenStudy (anonymous):

so you can make sure that you will be able to find the distance

OpenStudy (anonymous):

Definitely use the distance formula first

OpenStudy (anonymous):

Yeah, round to the nearest tenth to get a approximate answer

OpenStudy (anonymous):

Ok but how do I do that?

OpenStudy (anonymous):

http://www.purplemath.com/modules/distform.htm

OpenStudy (anonymous):

Use the distance formula, of course

OpenStudy (anonymous):

Google rocks!!!

OpenStudy (anonymous):

from -2 to -7 is 5 units, from -4 to 0 is 4 units, pythagoras says \(5^2+4^2=d^2\) so \(d=\sqrt{5^2+4^2}\)

OpenStudy (anonymous):

A(-7, -4) and B(-2, 0). D=sqrt(dx^2+dy^2) where dy and dx meand difference in x and y

OpenStudy (campbell_st):

the formula is \[d = \sqrt{(x _{2}-x _{1})^2 + ( y _{2}- y _{1})^2}\] designate the ordered pair A as the 1 point and the other B as the 2 point substitute and evaluate as shown above